OpenStudy (anonymous):

Write an equation in slope-intercept form of the line with the given slope and point: slope = 3 and (1, -2).

OpenStudy (anonymous):

explain please

OpenStudy (anonymous):

Do you know point slope form

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Use it here, in simplifying you will get into slope-intercept form

OpenStudy (anonymous):

how though? can you please help. they only gave me one set

OpenStudy (anonymous):

Well you know point slope form y-y1=m(x-x1) (x1,y1) is the point and m is the slope, so plug things in y+2=3(x-1) now you distribute the 3 on the right hand side and then isolate y